OpenStudy (anonymous):

HELP! I fan and give medals! Write the direct linear variation equation for the relationship: y varies directly with x and y = 4 when x = 12. A. y=1/3x B.y = x + 8 C.y = 3x D.y = x – 8

OpenStudy (anonymous):

@whpalmer4 @ParthKohli Help!! (Please) ;)

OpenStudy (anonymous):

I think it might be A. but I'm not sure. :)

OpenStudy (ranga):

Direct linear variation means: y = kx when x = 12, y = 4. Put the values above and solve for k.

OpenStudy (anonymous):

4=k times 12?

OpenStudy (ranga):

Yes. And therefore k = 4/12 = ?

OpenStudy (anonymous):

right :)

OpenStudy (anonymous):

0.333333

OpenStudy (ranga):

k = 1/3 and so your earlier answer was correct.

OpenStudy (anonymous):

oh ok thanks!!! :)

OpenStudy (ranga):

You are welcome.

OpenStudy (whpalmer4):

Yep, agreed that A is correct.